# Structural transitions in interacting lattice systems

### v1

@@ 1 Introduction
where ||$\cdot$|| ⋅ | is the Euclidean norm on $\mathbb{R}^{2}$, one can ask the following important questions:
@@ 2 Preliminary formalism
We say that $f\in\mathcal{F}$ if $f:(0,\infty)\to\mathbb{R}$, $|f(r^{2})|=O(r^{-2-\varepsilon})$ for some $\varepsilon>0$ as $r\to\infty$, and there exists a Radon measure $\mu_{f}$ on $(0,+\infty)$ such that
@@
where ||$\cdot$|| ⋅ | is the euclidean norm on $\mathbb{R}^{2}$.
@@
We use Gradshteyn-Ryzhik [18] notation for the third Jacobi theta function. Furthermore the subtraction of $-1$ in the square bracket is due to the absence of the self-energy term $(j,k)=(0,0)$ in the sum (2.4). If $\mu_{f}$ is absolutely continuous with respect to the Lebesgue measure, i.e. $d\mu_{f}(t)=\rho_{f}(t)dt$, then we have
@@
If the potential $f$ decays to zero at large distances too slowly, the integral in (2.5) may diverge which requires a regularization (analytic continuation). Like for instance, for the Riesz potential the integral converges if $s>2$. The Coulomb case with $s=1$ has to be regularized by adding a uniform neutralizing background which corresponds to the subtraction of the singular term $-\pi/t$ in the square bracket in the integral (2.5), as it is explained for instance in [32].
@@ 3 General theory of structural transitions
We use the fact that $f$ is analytic and $f(\cdot^{2})$ is absolutely summable at infinity. Therefore, the wished expansion easily yields from the one of the theta functions product, i.e. for all $t>0$ and as $\varepsilon\to 0$, the quantity $\theta_{3}\left({\rm e}^{-t\exp(-\varepsilon)}\right)\theta_{3}\left({\rm e}^{-t\exp(\varepsilon)}\right)$ is given by
@@
[…]
@@
    if $\+𝑓\subscriptf\in\mathcal{F}\backslash\mathcal{F}_{+}$ such that $\mu_{f}<0$ on $(0,r_{0})$ or on $(r_{1},+\infty)$ for some $r_{0},r_{1}>0$, then a transition point exists.
@@
  • if $A^{*}<A<A_{1}$, a minimizer $\varepsilon$ of $\varepsilon\to E(A,e^{\varepsilon})$ satisfies the following asymptotics, as $A\to A^{*}$:
@@
Since a transition point exists, a simple Taylor expansion of $E_{2}$ as $A\to A^{*}$ reads
@@
The value of $\varepsilon$ which minimizes $E(A,\cdot)$ satisfies the following asymptotic stationarity condition we get from (3.3):
@@
Furthermore, the singular behaviour $\varepsilon\propto(A-A^{*})^{\beta}$ as $A\to A^{*}$, $A>A^{*}$, of the minimizer defines the critical exponent $\beta$ which in our case acquires the mean-field value $1/2$. This critical exponent is universal in the sense that it does not depend on the particular form of the interaction potential $f$ as long as the (general) assumptions are satisfied.
@@
Therefore, there exists $A_{2}>0$ such that for $A^{t}<A<A_{2}$, a minimizer of $\varepsilon\to E(A,e^{\varepsilon})$ satisfies the following asymptotics as $A\to A^{t}$:
@@
This follows from (3.10) and (3.11) expanded to order $5$ – coming from the Taylor expansion of the energy to order $6$ as $E(A,e^{\varepsilon})=E_{0}(A)+E_{2}(A)\varepsilon^{2}+E_{4}(A)\varepsilon^{4}+E_{6}(A)\varepsilon^{6}+o(\varepsilon^{8})$ – and following exactly the same lines as in the proof of Proposition 3.8. ∎
@@
For $\alpha=\alpha^{t}$, we therefore get the singular behavior $\varepsilon\propto(A-A^{t})^{1/4}$ as $A\to A^{t}$, $A>A^{t}$, so that the critical exponent at the tricritical point is $\beta^{t}=1/4$ which is again universal up to the assumptions.
@@ 4 Double Yukawa potential
The potential (4.1) is supposed to have an attractive minimum $f(r_{\min})<0$ at some $r=r_{\min}$. For simplicity, we set $r_{\min}=1$ and $f(1)=-1$, reducing in this way the number of parameters four by two. We keep the independent parameters $(v_{1},\kappa_{1})$ since
@@
[…]
@@
[…]
@@ 5 Yukawa-Coulomb potential
has a minimum at $r_{\min}=1$ and $f(1)=-1$ under conditions
@@
It is clear that $f\not\in\mathcal{F}$ since it is not integrable at infinity. Therefore, a regularization of the divergent lattice sum by a neutralizing background is necessary for the Coulomb term [32]. The energy then reads as
@@
The neutralizing background manifests itself as the addition of the singular term $-\pi/t$ in the last square bracket, which ensures the convergence of the integral. It is again straightforward to show that transitions point exists by direct application of Proposition 3.6 in the same way that we did in Lemma 4.1 for the $\kappa_{2}>0$ case.
Figure 6: The plot of the energy difference $E(\Delta)-E(1)$ versus $\Delta$ for the Yukawa parameter $\kappa_{1}=2.0365$, just below its tricritical value $\kappa_{1}^{t}=2.036517758847\ldots$ where a first-order transition between the square and rectangular lattices exists. Three values of the inverse particle density $A$ are considered: $A=2.79544356250$ when the square lattice with $\Delta=1$ prevails, the first-order transition value $A_{\rm trans}=2.795443562576$ where the energies of the square and rectangular lattices, separated by energy barriers, coincide and $A=2.795443562606$ where the rectangular case is dominant via a jump in $\Delta$.

### v2

@@ 1 Introduction
where $|\cdot|$ is the Euclidean norm on $\mathbb{R}^{2}$, one can ask the following important questions:
@@ 2 Preliminary formalism
We say that $f\in\mathcal{F}$ if $f:(0,\infty)\to\mathbb{R}$, $|f(r^{2})|=O(r^{-2-\varepsilon})$ for some $\varepsilon>0$ as $r\to\infty$, $f\in C^{\infty}(\mathbb{R}_{+}^{*})$ and there exists a Radon measure $\mu_{f}$ on $(0,+\infty)$ such that
@@
where $|\cdot|$ is the euclidean norm on $\mathbb{R}^{2}$.
@@
Remark 2.8.
Since $f$ is admissible, it follows immediatly that, by composition and absolute summability of the energy (2.4), $E\in C^{\infty}(\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*})$.
@@
We use the Gradshteyn-Ryzhik [18] notation for the third Jacobi theta function. Furthermore the subtraction of 1-1 in the square bracket is due to the absence of the self-energy term $(j,k)=(0,0)$ in the sum (2.4). If $\mu_{f}$ is absolutely continuous with respect to the Lebesgue measure, i.e. $d\mu_{f}(t)=\rho_{f}(t)dt$, then we have
@@
If the potential $f\not\in\mathcal{F}$ decays to zero at large distances too slowly – like for instance for the Riesz potential with $s<2$ – the integral in (2.5) may diverge which requires a regularization. Typical examples are available in the literature, in particular in [22, Section IV.2], where the renormalization for Riesz potential and logarithmic energy are presented and where it is showed that adding a uniform neutralizing background to the system yields to the analytic continuation of the Epstein zeta function. In the Coulomb case with $s=1$, it corresponds to the subtraction of the singular term $\pi/t$ in the square bracket in the integral (2.5), as it is also explained in [33]. Therefore, the energy for
$f(r)=\frac{1}{r}=\int_{0}^{\infty}e^{-r^{2}t}\frac{dt}{\sqrt{\pi}\sqrt{t}},$
has to be redefined, using a straightforward change of variable, as
$E(A,\Delta)=\frac{1}{2\sqrt{A\pi}}\int_{0}^{\infty}\left[\theta_{3}(e^{-t\Delta})\theta_{3}(e^{-t/\Delta})-1-\frac{\pi}{t}\right]\frac{dt}{\sqrt{t}}.$
@@ 3 General theory of structural transitions
We use the fact that $f\in C^{\infty}(0,\infty)$ and $f$ is absolutely summable at infinity which implies that $E(A,\cdot)\in C^{\infty}(0,\infty)$ (see Remark 2.8). Therefore, the wished expansion easily yields from the one of the theta functions product, i.e. for all $t>0$ and as $\varepsilon\to 0$, the quantity $\theta_{3}\left({\rm e}^{-t\exp(-\varepsilon)}\right)\theta_{3}\left({\rm e}^{-t\exp(\varepsilon)}\right)$ is given by
@@
Let $f\in\mathcal{F}\backslash\mathcal{F}_{+}$, then any $A^{*}$ such that
@@
    if $f\in\mathcal{F}\backslash\mathcal{F}_{+}$ such that $\mu_{f}<0$ on $(0,r_{0})$ or on $(r_{1},+\infty)$ for some $r_{0},r_{1}>0$, then a transition point exists.
@@
  • if $A^{*}<A<A_{1}$, any local minimizer $\varepsilon$ of $\varepsilon\to E(A,e^{\varepsilon})$ satisfies the following asymptotics, as $A\to A^{*}$:
@@
Since a transition point exists, a simple Taylor expansion of $E_{2}$ as $A\to A^{*}$ – ensured by the fact that $E(\cdot,\Delta)\in C^{\infty}(\mathbb{R}_{+}^{*})$ (see Remark 2.8)– reads
@@
The value of $\varepsilon$ which locally minimizes $E(A,\cdot)$ satisfies the following asymptotic stationarity condition we get from (3.3):
@@
Furthermore, the singular behaviour $\varepsilon\propto(A-A^{*})^{\beta}$ as $\to A^{*}, A>A$$A^{*}$, of the minimizer (and actually any local one) defines the critical exponent $\beta$ which in our case acquires the mean-field value $1/2$. This critical exponent is universal in the sense that it does not depend on the particular form of the interaction potential $f$ as long as the (general) assumptions are satisfied.
@@
Therefore, there exists $A_{2}>0$ such that for $A^{t}<A<A_{2}$, any local minimizer of $\varepsilon\to E(A,e^{\varepsilon})$ satisfies the following asymptotics as $A\to A^{t}$:
@@
This follows from (3.10) and (3.11) expanded to order 55 – coming from the Taylor expansion of the energy to order 66, since $E\in C^{\infty}(\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*})$, as $E(A,e^{\varepsilon})=E_{0}(A)+E_{2}(A)\varepsilon^{2}+E_{4}(A)\varepsilon^{4}+E_{6}(A)\varepsilon^{6}+o(\varepsilon^{8})$ – and following exactly the same lines as in the proof of Proposition 3.8. ∎
@@
For $\alpha=\alpha^{t}$, we therefore get the singular behavior $\varepsilon\propto(A-A^{t})^{1/4}$ as $\to A^{t}, A>A$$A^{t}$, so that the critical exponent at the tricritical point is $\beta^{t}=1/4$ which is again universal up to the assumptions.
@@ 4 Double Yukawa potential
The potential (4.1) is supposed to have an attractive minimum $f(r_{\min})<0$ at some $r=r_{\min}$. For simplicity, we set $r_{\min}=1$ and f(1)=1f(1)=-1, reducing in this way the number of parameters four by two. We keep the independent parameters $(v_{1},\kappa_{1})$ since
@@
    The critical curve $\mathcal{C}$ ends up at the tricritical point with the coordinates $A^{t}\approx 2.7163619942262467$ and $v_{1}^{t}:=v_{1}(A^{t})\approx 6.7951845011079$, denoted by the cross.
@@
Figure 3: The double Yukawa potential with the parameter $\kappa_{1}=2$, the corresponding tricritical point has coordinates $A^{t}=2.7163619942262467\ldots$ and $v_{1}^{t}=6.7951845011079\ldots$. The log-log plot of $\Delta-1$ versus $A-A^{t}$ is presented in the rectangular region $A>A^{t}$, numerical data are denoted by open circles interconnected via a dashed line. The fitting via $\Delta-1\propto(A-A^{*})^{\beta}$ yields the tricritical exponent $\beta^{t}\approx 0.253$.
As concerns the tricritical point for the double Yukawa parameter $\kappa_{1}=2$, see the cross in Fig. 1, its coordinates $A^{t}=2.7163619942262467\ldots$ and $v_{1}^{t}=6.7951845011079\ldots$ were calculated by using the formula (3.14). The numerical calculation of the deviation $\Delta-1$ in the region of the rectangular lattice $A>A_{t}$ close to the tricritical point was made. The log-log plot of numerical data in Fig. 3 can be fitted as $\Delta-1\propto(A-A^{t})^{\beta^{t}}$, the obtained exponent $\beta^{t}\approx 0.253$ is reasonably close to the expected mean-field value $\beta^{t}=1/4$.
@@ 5 Yukawa-Coulomb potential
has a minimum at $r_{\min}=1$ and f(1)=1f(1)=-1 under conditions
@@
It is clear that $f\not\in\mathcal{F}$ since it is not integrable at infinity. Therefore, a regularization of the divergent lattice sum by a neutralizing background is necessary for the Coulomb term (see [33] and Remark 2.11). The energy then reads as
@@
The neutralizing background manifests itself as the addition of the singular term $\pi/t$ in the last square bracket, which ensures the convergence of the integral. It is again straightforward to show that transitions point exists by direct application of Proposition 3.6 in the same way that we did in Lemma 4.1 for the $\kappa_{2}>0$ case.
Figure 6: The plot of the energy difference $E(\Delta)-E(1)$ versus $\Delta$ for the Yukawa parameter $\kappa_{1}=2.0365$, just below its tricritical value $\kappa_{1}^{t}=2.036517758847\ldots$ where a first-order transition between the square and rectangular lattices exists. Three values of the inverse particle density AA are considered: $A=2.79544356250$ when the square lattice with $\Delta=1$ prevails, the first-order transition value $A_{\rm trans}=2.795443562576$ where the energies of the square and rectangular lattices, separated by energy barriers, coincide and $A=2.795443562606$ where the rectangular case is dominant via a jump in $\Delta$.